TUNABLE-INFRARED LASER DESORPTION OF THIN FILMS IN A TIME-OF-FLIGHT MASS SPEC-TROMETER

\underline{TIMOTHY CHENG}, MICHAEL A. DUNCAN, DEPARTMENT OF CHEMISTRY, UNIVERSITY OF GEORGIA, ATHENS GA, 30605.

Tunable-infrared laser desorption in a time-of-flight mass spectrometer has been shown to be a viable technique for probing thin films. This technique has been used on a number of thin films, such as coronene and C_{60}. By tuning the infrared laser, resonance enhnacement could be seen from the change of signal intensity depending on the wavelength. The spectrum and possible mechanisms for this procedure will be discussed.